# Bridging the Sim2Real Gap: Vision Encoder Pre-Training for Visuomotor Policy Transfer

### v1

@@ Abstract
Simulation offers a scalable and efficient alternative to real-world data collection for learning visuomotor robotic policies. However, the simulation-to-reality, or “Sim2Real” distribution shift—introduced by employing simulation-trained policies in real-world environments—frequently prevents successful policy transfer. This study explores the potential of using large-scale pre-training of vision encoders to address the Sim2Real gap. We examine a diverse collection of encoders, evaluating their ability to (1) extract features necessary for robot control while (2) remaining invariant to task-irrelevant environmental variations. We quantitatively measure the encoder’s feature extraction capabilities through linear probing and its domain invariance by computing distances between simulation and real-world embedding centroids. Additional qualitative insights are provided through t-SNE plots and GradCAM saliency maps. Findings suggest that encoders pre-trained on manipulation-specific datasets generally outperform those trained on generic datasets in bridging the Sim2Real gap. https://github.com/yyardi/Bridging-the-Sim2Real-Gap
@@ I Introduction
Robots that can reliably function in unstructured environments, like a home or construction site, operating from vision alone, have the potential to transform daily life and industry [1]. Visuomotor robotic policies—algorithms that enable robots to perform tasks solely using image observations are necessary to accomplish this. Recent advancements have been made using Neural Networks (NNs) and machine learning to learn these policies from data [2, 3, 4]. However, gathering the vast data necessary for robots to learn robust policies on physical robots is often impractical and can pose safety risks.
Simulation is a scalable alternative that has allowed researchers to replicate years of robot training in a fraction of the time [5]. Despite the incredible benefits, the simulation will inevitably differ from the real environment. This difference introduces a distribution shift known as the Sim2Real gap, which, if not adequately addressed, results in poorly performing simulation-trained policies when deployed in real settings.
A promising way to improve the Sim2Real gap is to pre-train visual representations [6]. These neural networks, trained on enormous datasets, convert raw images into compact, information-rich representation vectors that improve transferability by preferably extracting task-relevant information. This study assesses various encoders’ potential for transfer learning of visuomotor policies. We hypothesize that effective vision encoders capable of bridging the Sim2Real gap should (1) pick out relevant features from an image to infer actions (e.g., the objects and their position) and (2) ignore task-irrelevant parts of the image that change between simulated and real environments (e.g., the table color and clutter objects).
To comprehensively understand the potential for the vision encoders to (1) predict the correct actions based on the image and (2) to neglect domain-specific features that vary between simulated and real environments, we employ various quantitative and qualitative metrics described in detail in Section 3.
We find that the type of pre-training data used, either manipulation-specific data or general image data, tends to be the most important factor in determining an encoder’s overall performance. We also found that CNN-based encoders generally outperformed ViT-based encoders in domain invariance. We present our findings as a 2-dimensional plot according to our two criteria.
@@ II Related Works
Given the significant potential of AI-enabled robots capable of real-world unsupervised operation, many researchers have explored the nature of and potential solutions to the Sim2Real gap. Training machine learning models to help robots learn visuomotor policies traditionally requires enormous datasets that are costly and time-consuming to produce. Thus, it is crucial to solve the problem of increasing the viability of robots with diversely applicable visuomotor policies in the future.
@@ II-A Pre-Trained Visual Representations
The robotics research community has investigated pre-trained visual representations (PVRs) for training downstream policies in real-world tasks, comparing their performance across simulated and real environments. PVRs produce embeddings (output features of 1-dimensional vectors) extracted from visual data using pre-trained vision encoders designed for visuomotor tasks. Silwal et al. conducted a comprehensive investigation into using PVRs for training policies, evaluating their effectiveness in simulated and real environments [7]. The study demonstrated that PVRs significantly improved Sim2Real transfer in the ImageNav task compared to previous methods.
Burns et al. further studied the success of PVRs in robotic manipulation tasks, focusing on the attributes of pre-trained vision models robust to distribution shifts, including subtle lighting and texture changes [8]. The study evaluates fifteen models to identify which features made them more reliable under these conditions. The authors found that models pre-trained on manipulation-relevant data could consistently generalize better than models trained on standard pre-training datasets like ImageNet [9]. Notably, they found that a model’s strong emergent segmentation ability—the capacity to separate essential parts of an image from irrelevant details naturally—showcased its consistent success in generalization despite distribution shifts. Given the significant relevance of these distribution shifts to the Sim2Real gap, this finding applies directly to this inquiry’s application of the PVRs in examining whether the nature of the model’s pre-training indicates its performance.
@@ II-B Custom Frameworks for Learning to Generalize
II-C Probing to Evaluate Quality of Learning
The robotics community continues to explore cost-efficient methods for evaluating learned representations in training visuomotor policies. Linear probing is an affordable alternative to experimental assessments with reinforcement learning (RL) algorithms introduced in [11]. Researchers applied this approach in the Arcade Learning Environment, testing nine Atari games by comparing actions predicted by the probes to those taken by RL algorithms. The probes estimated task-specific rewards and identified optimal actions, with results showing a strong correlation between probe predictions and RL performance. This method proved to be up to 600 times more computationally efficient than RL-based evaluations. These findings validate this research’s novel probing to predict the action-inference performance of surveyed models without requiring direct experimentation on robots.
@@ III Methods
We evaluate several pre-trained vision encoders’ ability to learn effective visuomotor policies transferable to real-world applications. We posit that to tackle the Sim2Real gap, effective encoders should exhibit high action-inference accuracy—indicating relevant feature extraction ability—and low source determination ability, or high domain invariance, thus minimizing the Sim2Real gap. Encoder attributes are assessed by generating embeddings, using linear probes and distance metrics for quantitative comparison, and performing qualitative analyses with t-SNE plots and Grad-CAM saliency maps on image samples [12, 13]. With matching action and domain labels, the image dataset includes more than 50,000 frames of a robot arm trying to construct a small table—a challenging task—from simulated and real—world settings [14, 15]. This large collection offers a solid basis for assessing encoder generalization across domains because it contains a variety of situations, such as different illumination and object layouts.
@@ 18 Swin Transformer Transformer 1024 87 Self-supervised General
20 VGG16 CNN 4096 138 Supervised General
21 VGG19 CNN 4096 143 Supervised General
@@ III-A The Encoders
Twenty-three encoders are analyzed across all procedures to evaluate the performance of various model architectures in the Sim2Real gap, making this study contain, to our knowledge, the largest number of models examined in any study within this field [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Using action probing (Sec. 3.D), centroid distance (Sec. 3.C), t-SNE [12] (Sec. 3.E), and GradCAM [13] (Sec. 3.F), we explore how well models can discover relevant features for predicting visuomotor policies and how well models generalize their learning. Our encoder selection provides a broad and representative sampling of contemporary visual representation learning approaches, spanning multiple neural network structures, training methodologies, and datasets useful for training more robust models. Each encoder is categorized into Convolutional Neural Networks (CNNs) [30] and Vision Transformers (ViTs) [31]. CNNs are traditional approaches to image analysis designed to learn and extract features from images hierarchically. Alternatively, ViTs treat inputs like puzzles, relating pieces to create a holistic view of the image. We also compare models pre-trained on different data sets, some using manipulation datasets like Ego4D [32], and others using general data sets like ImageNet [9], to see if this pre-training data impacts Sim2Real performance. A complete list of the encoders is provided in Table 1, including architectural type, training approach, and size.
@@ III-B Generating Embeddings (Passing Images with Neural Networks)
We use each encoder to generate and store embeddings for all sim and real-world images. Embeddings are vector representations of the observed images the encoders generate, abstractly demonstrating learned representations and extracted features from the image. When fed an image, the neural networks convert the data into abstract vectors representing the image. These numbers are then passed through the layers of the neural networks that compress the data into smaller dimensions and perform calculations on it. The embeddings can be passed through an action prediction model to predict actions and learn visuomotor policies. Each pre-trained vision encoder will process images differently due to fundamental differences in the neural network designs, processing methods, and learned weights, demonstrating varying performance in our thorough qualitative and quantitative procedures. All model embeddings are normalized to ensure analyses are on the same scale across all benchmarks [33].
III-C Quantitatively Evaluating Domain Invariance With Centroid Distance Metric
We quantify each encoder’s domain invariance using a centroid distance metric, which measures the average difference between encoder-learned representations of pictures from simulated and actual environments to compare the extent of the Sim2Real gap. This approach (1) uses Principal Component Analysis (PCA) [34] to project all encoders into the same embedding space despite the encoders having different-size embeddings out of the box. Then, (2) min-max normalization is used to scale embeddings between 0 and 1, ensuring that scale differences are accounted for. Min-max normalization subtracts each feature’s minimal value and divides it by its range. (3) The embeddings are then divided into sim and real subsets, and the centroids are computed by finding the mean value for each embedding dimension for each subset. The final step (4) is the Euclidean distance calculation between sim and real centroids, measuring how closely the embeddings align. The outputted number subtracted from 1 is the “domain invariance score,” (DIS) calculated in equation (1)
$DIS=1-\left|\frac{1}{n_{r}}\sum_{i=1}^{n_{r}}z_{i}-\frac{1}{n_{s}}\sum_{i=1}^{n_{s}}z_{j}\right|_{2}^{2}$ (1)
Here, $\frac{1}{n_{r}}\sum_{i=1}^{n_{r}}z_{i}\in R^{d_{x}}$ is a vector of the same size as the embeddings produced by the encoder that represents the centroid of the embeddings produced from the images from the real domain, and similarly for the images from the simulated domain for the second term in the norm.
III-D Quantitatively Evaluating Task-Relevant Information With “Action Probing”
We quantitatively analyze each encoder’s ability to infer actions with a technique we call “action probing,” based on the interpretation technique of linear probing [11]. For any given encoder to succeed in a visuomotor policy learning application, it must create embeddings with features relevant to determine what motor commands the robot should follow. The probe is a simple single-layer neural network that maps the embeddings to a prediction of the motor commands the robot should take to complete a task. The process consists of two parts: (1) training, in which we measure loss from the probe against labeled data and adjust weights accordingly for the next epoch, and (2) validation, in which the model predicts an action from held-out images to ensure the model is not purely memorizing the image features when adjusting weights.
Figure 2: An Example Action Probe Training Process (ResNet34).
The linear probes were run on the embeddings of each encoder over 10,000 epochs, measuring both training and validation loss, and an average of the training loss over the last 100 epochs was taken. This average loss is subtracted from one to give the “action score,” which returns a comparative, quantitative metric of the accuracy of potential motor commands from the embeddings. The action score (AS) is calculated in equation (2).
$AS=1-\frac{1}{n}\frac{1}{d}\sum_{i=1}^{n}\sum_{j=1}^{d}\left|a_{ij}-g_{\phi}(z_{i})_{j}\right|^{2}$ (2)
In (2), $n$ is the number of data points across sim and real, $d$ is the dimensionality of each robot action, $a\in R^{n\times d}$ is the dataset of all the ground truth actions, $g_{phi}:R^{d_{x}}\xrightarrow{}R^{d}$ is the action probe function that maps the embedding vector $z_{i}$ of dimension $d_{z}$ for each image into a prediction of the corresponding action with parameters $\phi$ learned using gradient ascent on the above score function.
III-E Qualitatively Analyzing Domain Invariance with t-SNE Plotting
We use t-distributed Stochastic Neighbor Embedding (t-SNE) [12] plots to qualitatively analyze the differences between simulation-generated embeddings and real-world embeddings for each encoder based on the overlap of the plotted clusters. t-SNE is an unsupervised non-linear dimensionality reduction technique for data exploration and visualizing high-dimensional data. t-SNE-reduced clusters of the sim and real embeddings are plotted in 2-dimensions, with the presumption that less overlap between the clusters indicates encoders picking up on more spurious differences between the images. Increased distance would lead to different embeddings and, potentially, a more significant Sim2Real gap.
Figure 3: An Example t-SNE Plot (ResNet34)
III-F Qualitatively Understanding Encoders with GradCAM Saliency Maps
To comprehensively understand each model’s performance regarding the Sim2Real gap and their applicability in visuomotor policies, we qualitatively analyze what each model is “looking for” when given an image. As the images are converted into pixel values and passed through the model layers, the pre-trained weights will adjust them according to the target for which they have been trained (in this case, visuomotor control). GradCAM (Gradient-weighted Class Activation Mapping) is suitable for identifying the most critical regions of each input image for the learned representations [13]. These visual explanations are later compared with quantitative metrics to analyze each encoder’s action inference and domain invariance capabilities. GradCAM works by calculating a gradient of the activations in the final layer of a neural network and then overlaying this gradient onto the input image, representing the sensitivity of different regions to altered activations in the final layer as colors. The heat map indicates which parts of the image have the most significant impact on the generated embedding, thus demonstrating what the model is paying attention to. Some characteristics we can observe are whether the model focuses on the robot or the surrounding environment (which could impact domain-invariance) and how narrow the focus of the robot is on the robot’s end-effector and the objects on the table it interacts with (impacting action-inference).
Figure 4: An Example GradCAM Plot for the ResNet34 encoder, comparing visually similar images for sim (top row) and real (bottom row).

### v2

# Bridging the Sim2Real Gap: A Framework for Evaluating Vision Encoder Pre-Training for Visuomotor Policy Transfer

@@ Abstract
Simulation offers a scalable and efficient alternative to real-world data collection for learning visuomotor robotic policies. However, the simulation-to-reality, or “Sim2Real” distribution shift—introduced by employing simulation-trained policies in real-world environments—frequently prevents successful policy transfer. We present an offline framework to evaluate the performance of using large-scale pre-trained vision encoders to address the Sim2Real gap. We examine a diverse collection of encoders, assessing their ability to extract features necessary for robot control (Action Score) while remaining invariant to task-irrelevant environmental variations (Domain Invariance Score). Evaluating 23 encoders, we reveal patterns across architectures, pre-training datasets, and parameter scales. Our findings show that manipulation-pretrained encoders consistently achieve higher Action Scores, CNN-based encoders demonstrate stronger domain invariance than ViTs, and the best-performing models combine both properties, underscoring DIS and AS as complementary predictors of Sim2Real transferability.
Index Terms:
Sim2Real, vision encoder, transfer learning, robotic manipulation, feature extraction, domain invariance
@@ I Introduction
Robots that can reliably function in unstructured environments, operating from vision alone, have the potential to transform daily life and industry [1]. Visuomotor robotic policies—algorithms enabling robots to perform tasks solely using image observations—are necessary for this. Recent advancements have been made using Neural Networks (NNs) and machine learning to learn these policies from data [2, 3, 4]. However, gathering the vast data necessary for robots to learn robust policies on physical robots is often impractical and unsafe.
Simulation is a scalable alternative that allows researchers to rapidly replicate years of robot training [5]. Despite the incredible benefits, a simulated environment will inevitably differ from its real-world counterpart. This difference introduces a distribution shift known as the Sim2Real gap, which, if not adequately addressed, results in poorly performing simulation-trained policies when deployed in real environments.
Prior work has demonstrated that pre-trained vision encoders improve Sim2Real transfer, however, a central challenge remains: how can we evaluate representation quality offline, before expensive robot rollouts [6]. These neural networks, pre-trained on enormous datasets, convert raw images into compact, information-rich representation vectors that improve transferability by preferably extracting task-relevant information. This study assesses various encoders’ potential for transfer learning of visuomotor policies. We hypothesize that effective vision encoders capable of bridging the Sim2Real gap should pick out relevant features from an image to infer actions (e.g., the objects and their position) and ignore task-irrelevant parts of the image that change between simulated and real environments (e.g., the table color and clutter objects).
In this paper, we propose a framework for offline evaluation of encoders in Sim2Real contexts. The framework is built around two quantitative metrics: the Domain Invariance Score (DIS), which measures the alignment of simulated and real-world embeddings, and the Action Score (AS), which evaluates how well encoder embeddings capture task-relevant information through linear probing. We incorporate Grad-CAM saliency maps as a qualitative validation tool, providing interpretable visual evidence that high-performing encoders attend to task-relevant features across both simulated and real domains.
@@ II Related Works
Given the significant potential of AI-enabled robots capable of real-world unsupervised operation, studies have explored the nature of the Sim2Real gap and devised potential solutions.
@@ II-A Pre-Trained Visual Representations
The robotics research community has investigated pre-trained visual representations (PVRs) for training downstream policies in real-world tasks, comparing their performance across simulated and real environments [7]. PVRs produce embeddings (output features of 1-dimensional vectors) extracted from visual data using pre-trained vision encoders designed for visuomotor tasks. Silwal et al. [7] conducted a comprehensive investigation into using PVRs for training policies, evaluating their effectiveness in simulated and real environments. The study demonstrated that PVRs significantly improved Sim2Real transfer in the ImageNav task compared to previous methods.
Burns et al. [8] further studied the success of PVRs in robotic manipulation tasks, focusing on the attributes of pre-trained vision models robust to distribution shifts, including subtle lighting and texture changes. The study evaluates fifteen models to identify which features made them more reliable under these conditions. The authors found that models pre-trained on manipulation-relevant data could consistently generalize better than models trained on standard pre-training datasets like ImageNet [9]. Notably, they found that a model’s strong emergent segmentation ability—the capacity to separate essential parts of an image from irrelevant details naturally—showcased its consistent success in generalization despite distribution shifts. Given the significant relevance of these distribution shifts to the Sim2Real gap, this finding applies directly to this inquiry’s application of the PVRs in examining whether the nature of the model’s pre-training indicates its performance.
@@ III Methods
We develop a framework to evaluate several pre-trained vision encoders’ ability to learn effective visuomotor policies transferable to real-world applications. We posit that to tackle the Sim2Real gap, effective encoders should exhibit high action-inference accuracy—indicating relevant feature extraction ability—and low source determination ability, or high domain invariance, measured through the Domain Invariance Score (DIS). Encoder attributes are assessed by generating embeddings, using our framework of linear probes and distance metrics for quantitative comparison using the Action Score (AS) [11].
With matching action and domain labels, the image dataset utilized more than 50,000 frames of a robot arm trying to construct a small table from simulated and real-world environments [12]. This large collection offers a solid basis for assessing encoder generalization across domain shifts, including different illumination and object layouts. [13]
@@ 18 Swin Transformer Transformer 1024 87 Self-supervised General
20 VGG-16 CNN 4096 138 Supervised General
21 VGG-19 CNN 4096 143 Supervised General
@@ III-A The Encoders
Twenty-three encoders are analyzed across all metrics to evaluate the performance of various model architectures in the Sim2Real gap, making this study contain, to our knowledge, the largest number of models examined in any study within this field [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. Our encoder selection provides a broad and representative sampling of contemporary visual representation learning approaches, spanning multiple neural network structures, training methodologies, and datasets useful for training more robust models. Each encoder is categorized into Convolutional Neural Networks (CNNs) [28] and Vision Transformers (ViTs) [29]. CNNs are traditional approaches to image analysis designed to learn and extract features from images hierarchically. Alternatively, ViTs treat inputs like puzzles, relating pieces to create a holistic view of the image. We also compare models pre-trained on different data sets, some using manipulation datasets like Ego4D, and others using general data sets like ImageNet [9], to see if this pre-training data impacts Sim2Real performance. A complete list of the encoders is provided in Table I, including architectural type, training approach, and size.
@@ III-B Generating Embeddings (Passing Images with Neural Networks)
We use each encoder to generate and store embeddings for all sim and real-world images. Embeddings are vector representations of the observed images the encoders generate, abstractly demonstrating learned representations and extracted features from the image. Each pre-trained vision encoder will process images differently due to fundamental differences in the neural network designs, processing methods, and learned weights, demonstrating varying performance in our thorough qualitative and quantitative procedures. All model embeddings are normalized to ensure analyses are on the same scale across all benchmarks.
III-C Domain Invariance Score (DIS) Using Centroid Distances
With our framework, we quantify each encoder’s domain invariance using a centroid distance metric, which measures the average difference between generated embeddings of simulated and real images to compare the extent of the Sim2Real gap across models.
Let $f_{\theta}:\mathbb{R}^{H\times W\times C}\!\to\!\mathbb{R}^{d}$ be an encoder and $z=f_{\theta}(x)$ its embedding. To compare encoders with differing output dimensions on a common space, we concatenate all embeddings (sim+real) into $E\in\mathbb{R}^{n\times d}$, standardize features, and project to a fixed dimension $d^{\star}$ via PCA: $\tilde{E}=\mathrm{PCA}_{d^{\star}}(\mathrm{Std}(E))$. We then apply per-feature min–max normalization to obtain $\hat{E}$ with components in $[0,1]$. Denote simulated and real subsets by $\hat{E}_{\mathrm{sim}}$ and $\hat{E}_{\mathrm{real}}$, and their centroids
$\mu_{\mathrm{sim}}=\frac{1}{n_{s}}\sum_{j=1}^{n_{s}}\hat{z}_{j}^{\mathrm{sim}},\qquad\mu_{\mathrm{real}}=\frac{1}{n_{r}}\sum_{i=1}^{n_{r}}\hat{z}_{i}^{\mathrm{real}}.$
The Domain Invariance Score (DIS) is the inverse, dimension-normalized Euclidean gap between centroids,
$\mathrm{DIS}\;=\;\Big{[}\,1-\frac{\lVert\mu_{\mathrm{real}}-\mu_{\mathrm{sim}}\rVert_{2}}{\sqrt{d^{\star}}}\,\Big{]}_{+},$ (1)
with $t]_{+}=\max\{0,t\}$. This choice follows two-sample testing intuition: with a linear kernel, domain discrepancy reduces to the distance between mean embeddings; thus a smaller $\mathrm{DIS}$ on 4 indicates stronger sim–real alignment.
Input: Embeddings $E\in\mathbb{R}^{n\times d}$, flags $D\in\{\mathrm{sim},\mathrm{real}\}^{n}$, target dim $d^{\star}$
$\tilde{E}\leftarrow\mathrm{PCA}_{d^{\star}}(\mathrm{Std}(E))$ ;
  // standardize then project
for $j=1$ to $d^{\star}$ do
 $m_{j}\leftarrow\min_{i}\tilde{E}_{ij}$, $M_{j}\leftarrow\max_{i}\tilde{E}_{ij}$, $\hat{E}_{ij}\leftarrow\frac{\tilde{E}_{ij}-m_{j}}{\max(M_{j}-m_{j},\;\varepsilon)}$
Partition $\hat{E}$ into $\hat{E}_{\mathrm{sim}}$ and $\hat{E}_{\mathrm{real}}$ using $D$;
$\mu_{\mathrm{sim}}\leftarrow\frac{1}{|\hat{E}_{\mathrm{sim}}|}\sum_{z\in\hat{E}_{\mathrm{sim}}}z$;
$\mu_{\mathrm{real}}\leftarrow\frac{1}{|\hat{E}_{\mathrm{real}}|}\sum_{z\in\hat{E}_{\mathrm{real}}}z$;
$d\leftarrow\lVert\mu_{\mathrm{real}}-\mu_{\mathrm{sim}}\rVert_{2}\big{/}\sqrt{d^{\star}}$;
return $\mathrm{DIS}\leftarrow\max(0,\,1-d)$;
Algorithm 1 Domain Invariance Score (DIS) with PCA + Min–Max + Euclidean
III-D Action Score (AS) Using Probing
Our framework can quantitatively analyze each encoder’s ability to infer actions with a technique we call “action probing,” based on the interpretation technique of linear probing [11].
Given frozen encoder embeddings $z_{i}\in\mathbb{R}^{d_{z}}$ and ground-truth robot actions $a_{i}\in\mathbb{R}^{d_{a}}$, we assess task-relevant information via a linear probe $g_{\phi}(z)=Wz+b$ trained with mean-squared error on a train split and evaluated on a held-out validation split. The probe solves
$\min_{\phi}\;\frac{1}{n_{\mathrm{tr}}}\sum_{i\in\mathcal{D}_{\mathrm{tr}}}\lVert a_{i}-g_{\phi}(z_{i})\rVert_{2}^{2},$
and we define the Action Score (AS) from the validation error as
$\mathrm{AS}\;=\;\Big{[}\,1-\frac{1}{n_{\mathrm{val}}\,d_{a}}\sum_{i\in\mathcal{D}_{\mathrm{val}}}\lVert a_{i}-g_{\phi}(z_{i})\rVert_{2}^{2}\,\Big{]}_{+},$ (2)
where $t]_{+}=\max\{0,t\}$. Thus, higher $\mathrm{AS}$ on 4 indicates that actions are linearly recoverable from the encoder’s representation (greater task informativeness).
Input: Train/val embeddings $\{(z_{i},a_{i})\}$ with $z_{i}\!\in\!\mathbb{R}^{d_{z}}$, $a_{i}\!\in\!\mathbb{R}^{d_{a}}$; epochs $T$, batch size $B$
Initialize linear probe $g_{\phi}(z)=Wz+b$;
for $t=1$ to $T$ do
 for minibatch $(Z,A)\subset\mathcal{D}_{\mathrm{tr}}$ of size $B$ do
    $\phi\leftarrow\phi-\eta\nabla_{\phi}\frac{1}{B}\sum\lVert A-g_{\phi}(Z)\rVert_{2}^{2}$
Compute validation MSE: $\mathrm{MSE}_{\mathrm{val}}\leftarrow\frac{1}{n_{\mathrm{val}}\,d_{a}}\sum_{i\in\mathcal{D}_{\mathrm{val}}}\lVert a_{i}-g_{\phi}(z_{i})\rVert_{2}^{2}$;
return $\mathrm{AS}\leftarrow\max(0,\,1-\mathrm{MSE}_{\mathrm{val}})$;
Algorithm 2 Action Score (AS) via Linear Probing
III-E Qualitatively Understanding Encoders with GradCAM Saliency Maps
Grad-CAM [30] is used to visualize encoder attention on robot scenes. We do not present this as a contribution, but as a supporting check: encoders with low DIS should attend similarly to task-relevant regions across sim and real images, while encoders with high AS should focus sharply on manipulators and objects. This ensures our quantitative metrics correspond to meaningful behaviors.